### v1

# Microquasar jet-cocoon systems as PeVatrons

Abstract
The origin of Galactic cosmic rays (CRs), particularly around the knee region ($\sim$3 PeV), remains a major unsolved question. Recent observations by LHAAASO suggest that the knee is shaped mainly by protons, with a transition to heavier elements at higher energies. Microquasars – compact jet-emitting sources – have emerged as possible PeV CR accelerators, especially after detections of ultrahigh-energy gamma rays from these systems. We propose that the observed proton spectrum (hard below a few PeV, steep beyond) arises from the reacceleration of sub-TeV Galactic CRs via shear acceleration in large-scale microquasar jet-cocoon structures. Our model also naturally explains the observed spectrum of energies around a few tens of PeV by summing up heavier nuclei contributions. Additionally, similar reacceleration processes in radio galaxies can contribute to ultrahigh-energy CRs, bridging Galactic and extragalactic origins. Combined with low-energy CRs from supernova remnants and galaxy clusters around the second knee region, this scenario could provide a unified explanation for CRs across the entire energy spectrum.
pacs:
Valid PACS appear here
preprint: APS/123-QED
I Introduction
The origin of cosmic rays (CRs), mainly composed of protons, heliums, and heavier nuclei, is a long-standing puzzle in astrophysics Antoni et al. (2005); Kotera and Olinto (2011); Kachelriess and Semikoz (2019). The energy spectrum of CRs spans over a wide range from $10^{9}\rm~{}eV$ to $10^{20}\rm~{}eV$, roughly following a power-law spectrum $\Phi_{\rm CR}\propto E^{-\gamma}$. The spectrum exhibits a distinctive feature with a change in the power law spectral index from $\gamma\sim 2.7$ to $\gamma\sim 3$ at the energy around $4\rm~{}PeV$, known as knee Antoni et al. (2005); Amenomori et al. (2008); Cao et al. (2024a), discovered by Kulikov and Khristiansen in 1958 Kulikov and Khristiansen . At energies around $\sim 100\rm~{}PeV$, the spectral index changed from $\gamma\sim 3$ to $\gamma\sim 3.3$, called the “second knee” (e.g., Nagano et al., 1992; Abbasi et al., 2018). The spectral index changes back to $\gamma\sim 2.5$ at the ankle around 4 EeV (e.g., Bird et al., 1993), and the flux is largely suppressed at a few tens of EeV Abbasi et al. (2008); Abraham et al. (2008). The knee is believed to be formed due to the limited maximum rigidity of Galactic CR accelerators Hillas (2005), or reflecting a spectral break at the same rigidity Cardillo et al. (2015). The knee could also be formed due to the change of the energy dependent diffusion coefficient in the propagation of CRs Giacinti et al. (2015), while CRs between the knee and ankle should be marked as the transition region from Galactic to extragalactic CRs Gaisser (2006). CRs with energy above EeV are regarded as ultrahigh-energy cosmic rays (UHECRs), which have an extragalactic origin as supported by the observation of the large-scale anisotropy, where the dipole direction points in the opposite direction of the Galactic center Aab et al. (2018).
CR composition is the key to identifying the origins. The general pattern of CR abundances around a few GeV per nucleon resembles that of the interstellar medium or primordial solar composition, with peaks in the CNO and Fe region Drury (2018). Recently, a knee-like feature around $3\rm~{}PeV$ is also revealed on the measured all-particle spectrum of CRs. The logarithm of the mean mass of CRs measured by the Large High Altitude Air Shower Observatory (LHAASO) showed trends toward a heavier composition below several hundred TeV and above several PeV, while it is a light composition dominated around the knee, and the turning point is consistent with the position of the knee at energy around $3\rm~{}PeV$ Cao et al. (2024a). In particular, LHAASO reported the first high-purity proton spectrum, implicating the knee at the energy around $3\rm~{}PeV$ is related to the change of the proton spectrum Cao et al. (2025). The observed hardening feature above several hundreds of TeV implicates the emergence of a new component that has the maximum energy at least up to a few PeV Cao et al. (2025).
The physical origin of the knee is still unknown and should be related to Galactic PeVtrons. SNRs are believed to be responsible for CRs at least up to the knee Reynolds (2008); Blasi (2013). However, SNRs as PeVatrons are challenged by both theoretical and observational studies Bell (2004, 2005); Gabici et al. (2019); Giuliani and Cardillo (2024), potentially implying existence of another type of PeVatrons. Potential candidates for PeVtrons include massive stars and stellar clusters Casse and Paul (1982); Cesarsky and Montmerle (1983); Aharonian et al. (2019), pulsars and pulsar winds Bednarek and Protheroe (2002); Giller and Lipski (2002); Bednarek and Bartosik (2004); Ohira et al. (2018), superbubbles Bykov et al. (2020), hypernovae Sveshnikova, L. G. (2003) and interacting supernovae Murase et al. (2014); Inoue et al. (2021), and isolated black holes Ioka et al. (2017); Kimura et al. (2025). The detection of very high-energy (VHE) gamma rays has shown that microquasars are promising sites for accelerating CRs around PeV and above, indicating potential sources of Galactic PeV CRs Cao et al. (2024b). Microquasars are X-ray binaries with relativistic jets Mirabel and Rodriguez (1994) and display phenomenology similar to active galactic nuclei (AGNs), although the size of the system is orders of magnitude different. It is suggested that shear and one-shot acceleration in the large-scale jet of radio galaxies can be responsible for the origin of UHECRs Kimura et al. (2018a); Caprioli (2015); Mbarek and Caprioli (2019); Mbarek et al. (2023, 2025). This motivates us to investigate whether the shear acceleration by microquasar jets can be Galactic PeVatrons. Previous studies also suggest microquasars as PeVatrons Cooper et al. (2020); Kimura et al. (2021); Yue et al. (2024), but these studies consider CR acceleration by shocks in jets or magnetic reconnections in accretion flows.
The paper is organized as follows. In Section II, we study the physics of Galactic microquasars and the shear acceleration process in the jet-cocoon system of microquasars. In Section III, we propagate the escaped CRs from microquasars and compare the results with observations. In Section IV, we combine CRs from Galactic SNRs and contributions from the shear acceleration process from extragalactic clusters of galaxies and radio galaxies, we proposed an overall picture of the origin of CRs on a wide energy range from GeV to $\sim 100\rm~{}EeV$. We give a summary and discussion in Sec. V.
II Cosmic rays accelerated in microquasars
II.1 The physics of microquasar jets
A microquasar consists of a compact object, usually a black hole, and a companion star, see Fig. 1 for a schematic plot. Currently, over 20 microquasars have been identified in our Galaxy Tetarenko et al. (2016); Corral-Santana et al. (2016). In the microquasar systems, material of the companion star falls to the compact star, which releases a large amount of the gravitational energy. A part of the gravitational energy is converted to outflow kinetic energy. These outflows, called jets, are launched at the vicinity of the compact star, narrowly collimated to a small solid angle, and often reach relativistic speeds. The kinetic power of the jets can reach the Eddington luminosity, $L_{j}\sim L_{\rm Edd}\sim 1\times 10^{39}(M/10M_{\odot})\rm~{}erg~{}s^{-1}$. The jets propagate in interstellar medium and extend to $\sim 100$ pc scales, which is many orders of magnitude larger than the scale of the compact star. See Ref. Done et al. (2007) for a comprehensive review.
Figure 1: A schematic plot of the reacceleration of Galactic low-energy GeV-TeV CRs in the large-scale jet-cocoon system of microquasars.
Recently, Ref. Cao et al. (2024b) reported detection of gamma rays above 100 TeV from five microquasars, three of which (SS 433, V4641 Sgr, and GRS 1915+105) exhibit extended gamma-ray morphology whose emission zone is as large as $\sim 50-100$ pc . These gamma-rays are most likely originating from the radiation of nonthermal particles accelerated in the extended jets Alfaro et al. (2017); Aharonian et al. (2024); Cao et al. (2024b). The other two (Cyg X-1 and MAXI J1820+070) are consistent with point source, possibly suggesting a compact emission region Fang et al. (2024); Kuze et al. (2025), although the best fit positions of gamma-ray sources have slight offsets from those of X-ray binaries.
SS 433 is a binary system consisting of a compact object and a supergiant star, with a pair of jets moving in opposite directions with launch velocity $v\sim 0.26\rm~{}c$, where $c$ is the speed of light Abell and Margon (1979). The extended two-sided jets with $\sim 40$ pc are detected in X-rays Safi-Harb and Ögelman (1997), and radio lobes surrounding the system extends up to $\sim 150$ pc Dubner et al. (1998); Sakemi et al. (2023). VHE gamma-ray emission from SS 433 has been detected at the jet lobes, which are spatially coincident with the X-ray hotspots Abeysekara et al. (2018); Aharonian et al. (2024); Cao et al. (2024b); Alfaro et al. (2024a). The observed gamma-ray emission is consistent with inverse Compton emission from a single population of electrons, and the magnetic field strength inside the jet is $\sim 20\rm~{}\mu G$. Refs. Sudoh et al. (2020); Kimura et al. (2020) pointed out that nonthermal electrons accelerated at SS 433 jets achieves PeV energies with a small Bohm factor, based on multiwavelength modeling. In addition, Ref. Ohmura et al. (2021) performed magnetohydrodynamics (MHD) simulations of the propagation of the jet in the surrounding medium, where the jet is spheroidal at early times and the wings are developed by the broadening spherical cocoon.
V4641 Sgr is another interesting microquasar who emits gamma-rays above 100 TeV. The morophorogy of the gamma-ray signal is clear two-sided lobe structure, and the size of the lobe is roughly $\sim 100$ pc Cao et al. (2024b); Alfaro et al. (2024b). Spatially extended X-ray signature is also detected from this objects, although the X-ray signal is concentrated on the central region within $\sim 15$ pc from the central black hole (Suzuki et al., 2025). The multiwavelength modeling of this system indicates $B\gtrsim 10\rm\mu G$ at the X-ray emission region.
GRS 1915+105 is a highly variable object that exhibits super-luminal motion with extended jets of sub-pc scales Fender et al. (1999). Its Eddington ratio used to be $\sim 0.1-1$ (Done et al., 2004) and suddenly decreased to 0.01 in 2018 Koljonen and Hovatta (2021). The gamma-ray source detected by LHAASO exhibits an extension of $\sim 50$ pc Cao et al. (2024b), which is consistent with the two-sided lobe structure detected by ALMA Tetarenko et al. (2018).
Nonthermal particles can be accelerated through the energy dissipation within the jet. With the luminosity requirement Blandford (2000) obtained from the Hillas condition (Hillas, 1984), the maximum possible energy of the accelerated CRs is estimated to be
$E_{A,\rm max}$ $\approx 3\times{10}^{16}~{}{\rm eV}~{}Z\epsilon_{B}^{1/2}$
×$\times{\left(\frac{L_{j}}{{10}^{39}~{}{\rm erg/s}}\right)}^{1/2}\left(\frac{\Gamma}{1.05}\right)^{-1}\left(\frac{\beta}{0.3}\right)^{1/2}$ (1)
where $Z$ is the nuclear charge, $\epsilon_{B}$ is the energy fraction carried by the magnetic field, $L_{j}$ is the total isotropic-equivalent luminosity of the jet, $\Gamma$ is the Lorentz factor, and $\beta$ is the characteristic velocity normalized by the speed of light. In our Galaxy, there are likely a few dozens of BH X-ray binaries with an X-ray luminosity of $L_{X}\gtrsim 10^{38}\rm~{}erg~{}s^{-1}$, each exhibiting sufficient luminosity, $L_{j}\sim 10^{39}\rm~{}erg~{}s^{-1}$, to potentially accelerate CRs up to the knee energy range. Here, we use frequently used assumptions of $L_{X}\sim 0.1\dot{M}c^{2}$ and $L_{j}\sim\dot{M}c^{2}$ to convert $L_{X}$ to $L_{j}$, where $\dot{M}$ is the mass accretion rate onto the BH.
II.2 Shear acceleration in microquasar jets
Shear acceleration, a class of the Fermi acceleration, has been suggested to be an efficient mechanism to accelerate CRs, where charged particles gain energy inside an ordered shear velocity field Berezhko and Krymskii (1981); Earl et al. (1988); Ostrowski (1990); Webb (1989); Rieger (2019). It has been shown that the discrete shear acceleration Ostrowski (1990) occurring on the surface of the jet-cocoon system in radio galaxies may accelerate low-energy “seed” galactic CRs to the UHE energy range, and the predicted spectrum and composition are consistent with observations Kimura et al. (2018a); Seo et al. (2024).
In AGNs, the reacceleration may occur through the “one-shot” acceleration, where the “seed” CRs receive a boost of a factor of $\Gamma^{2}$ Gallant and Achterberg (1999), and it is possible for powerful radio galaxies with a Lorentz factor of $\Gamma\gtrsim 30$ to accelerate “seed” galactic CRs with energy $E\lesssim 10^{17}\rm~{}eV$ to the UHE energy range Caprioli (2015); Mbarek and Caprioli (2019).
We consider the shear acceleration by cylindrical jets and cocoon as discussed in Kimura, Murase and Zhang (hereafter KMZ18) Kimura et al. (2018a). Based on discussion in Section II.1, we assume that the radius of the jet is $r_{\rm jet}=3\rm~{}pc$, the length is $l_{\rm jet}=100\rm~{}pc$, the jet velocity normalized by the speed of light is $\beta_{j}=0.3$ (corresponding to Lorentz factor $\Gamma_{j}\simeq 1.05$), the radius of the cocoon is $r_{\rm coc}=60\rm~{}pc$. The coherence length inside the cocoon is set to $l_{\rm coh}\sim 0.03r_{\rm coc}\sim 1.8\rm~{}pc$. The mean free path of CRs is given by $\lambda\sim l_{\rm coh}(E/E_{\rm coh})^{\delta}$, where $\delta=1/3$ for Kolmogorov turbulence and $E_{\rm coh}=ZeB_{\rm coc}l_{\rm coh}\sim 17Z(B_{\rm coc}/10{\rm~{}\mu G})(l_{\rm coh}/1.8\rm~{}pc)\rm~{}PeV$. CRs will undergo a discrete shear acceleration when the CR mean free path is longer than the width of the sheath layer, $\lambda\gtrsim r_{\rm sl}$. For $r_{\rm sl}\sim f_{\rm sl}r_{\rm jet}\sim 0.06(f_{\rm sl}/0.02)\rm~{}pc$, the injection energy is given by $\lambda\sim r_{\rm sl}$, which can be rewritten as
$E_{\rm inj}$$\approx$ $E_{\rm coh}\left(\frac{r_{\rm sl}}{l_{\rm coh}}\right)^{3}$
$\sim$ $0.6\left(\frac{B_{\rm coc}}{\rm 10\mu G}\right)\left(\frac{r_{\rm sl}}{0.06\rm~{}pc}\right)^{3}\left(\frac{l_{\rm coh}}{\rm 1.8pc}\right)^{-2}\rm~{}TeV.$
Thus, CRs with energies above several hundreds of GeV can undergo a discrete shear acceleration mechanism. We should note that the width of the shear layer and injection energy to the discrete shear acceleration are still uncertain, and $f_{\rm sl}\sim 1$ is suggested by recent high-resolution MHD simulations when the kink instability is fully developed Wang et al. (2024).
Based on KMZ18 Kimura et al. (2018a), by balancing the acceleration and confinement timescales, the peak energy in the energy spectrum of escaping CR protons can be estimated to be
$E_{\rm max}$ $=$ $\zeta eB_{\rm coc}l_{\rm coh}^{1/2}r_{\rm jet}^{1/2}\Gamma_{\rm jet}\beta_{\rm jet}$
$\simeq$ $4.3{\rm~{}PeV}\left(\frac{B_{\rm coc}}{10\mu\rm G}\right)\left(\frac{l_{\rm coh}}{1.8\rm~{}pc}\right)^{1/2}\left(\frac{r_{\rm jet}}{3\rm pc}\right)^{1/2}\left(\frac{\beta_{\rm jet}}{0.3}\right)\zeta,$
where $\zeta\equiv(\zeta_{c}/\zeta_{a})^{2}$ is a numerical factor of order unity.
The average rate of CRs injected into the shear acceleration is estimated to be
$\dot{N}_{\rm CR}$ $\approx\frac{N_{i,\rm inj}}{t_{\rm ad}}\approx\frac{2\pi r_{\rm jet}^{2}l_{\rm jet}n_{\rm cr}}{r_{\rm coc}/v_{\rm exp}}$
$\simeq 7.1\times 10^{33}{\rm~{}s^{-1}}\left(\frac{r_{\rm jet}}{3\rm~{}pc}\right)^{2}\left(\frac{l_{\rm jet}}{100\rm~{}pc}\right)$ (4)
×$\times\left(\frac{r_{\rm coc}}{60\rm~{}pc}\right)^{-1}\left(\frac{v_{\rm exp}}{3000\rm~{}km~{}s^{-1}}\right),$
where $t_{\rm ad}\approx r_{\rm coc}/v_{\rm exp}$ is the adiabatic cooling time, $v_{\rm exp}$ is the expansion velocity and $n_{\rm cr}\sim 2.2\times 10^{-14}\rm~{}cm^{-3}$ is the observed number density of CRs at rigidity 0.32 TV. Assuming that CRs are accelerated to PeV energies, the total luminosity of CRs accelerated by a microquasar is roughly estimated to be
$L_{\rm CR}$ $\approx E_{\rm max}\dot{N}_{\rm CR}$ (5)
$\simeq 5.0\times 10^{37}\left(\frac{r_{\rm jet}}{3\rm~{}pc}\right)^{5/2}\left(\frac{l_{\rm jet}}{100\rm~{}pc}\right)\left(\frac{v_{\rm exp}}{3000\rm~{}km~{}s^{-1}}\right)$
×$\times\left(\frac{r_{\rm coc}}{60\rm~{}pc}\right)^{-1}\left(\frac{B_{\rm coc}}{10\mu\rm G}\right)\left(\frac{l_{\rm coh}}{1.8\rm~{}pc}\right)^{1/2}\left(\frac{\beta_{\rm jet}}{0.3}\right)\zeta\rm~{}erg~{}s^{-1}.$
Parameters $r_{\rm jet}$ $B_{\rm jet}$ $\beta_{\rm jet}$ $r_{\rm coc}$ $B_{\rm coc}$ $l_{\rm coh}$ $l_{\rm jet}$
[pc] [$\rm\mu G$] [pc] [$\rm\mu G$] [pc] [pc]
Value 3 50 0.3 60 10 1.8 100
Table 1: Model parameters used in Monte Carlo simulations.
We perform Monte Carlo simulations with the code used in KMZ18 Kimura et al. (2018a) with parameters given in Table 1, and the resulting energy spectra for escaping CR protons are shown in Fig. 2. As seen in the figure, our shear acceleration scenario can achieve spectral peak around PeV energies with reasonable parameters (see Section V for discussion of parameter choices). The escaping CR proton spectrum show a hard spectrum below the peak and a steep power-law decay beyond the peak.
Figure 2: The energy spectrum of escaping CRs predicted in the shear acceleration mechanism from Monte Carlo simulations. The parameters are listed in Table 1. The above plot is multiplied with $\rm exp(-\mathcal{R}/\mathcal{R}_{\rm max})$, where $\mathcal{R}_{\rm max}=4\times 10^{7}\rm~{}GeV$ is the maximum acceleration energy from microquasar jets.
III Observed cosmic-ray spectrum on Earth
Figure 3: Predicted CR intensity on Earth and the mean mass $\langle{\rm ln}(A)\rangle$ distribution from microquasars. The all-particle energy spectra are taken from LHAASO Cao et al. (2024a), HAWC Alfaro et al. (2017), NUCLEON Grebenyuk et al. (2019), TIBET Amenomori et al. (2008), TA Matthews (2017), and Auger Verzi (2019), while the proton spectrum are taken from LHAASO Cao et al. (2025), KASCADE-Grande Arteaga-Velázquez et al. (2017), and IceTop Aartsen et al. (2019). The measurement of the mean mass $\langle{\rm ln}(A)\rangle$ distribution are taken from LHAASO Cao et al. (2024a) and IceTop Aartsen et al. (2019).
In general, the propagation of CRs in our Galaxy can be described by the transport equation Parker (1965); Gaisser et al. (2016),
$\frac{\partial N_{i}(E,\mathbf{x})}{\partial t}+\mathbf{V}\cdot\nabla N_{i}(E,\mathbf{x})-\nabla\cdot\left[D(r)\nabla N_{i}(E,\mathbf{x})\right]$
$=Q_{i}(E,\mathbf{x},t)-p_{i}N_{i}(E,\mathbf{x})$
[…]
where $N_{i}(E,\mathbf{x})$ is the number density per energy at position $\mathbf{x}$, the second and third term describes the time evolution of particle number density due to advection and spatial diffusion, respectively. The first term in the right-hand represents injection of particles of type $i$, and the second and last term accounts for the loss and gain processes. Here, the loss of nuclei of type $i$ caused by collisions and decay with a rate $p_{i}=v\rho/\lambda_{i}+1/\gamma\tau_{i}$, where $\rho$ is the interstellar medium density, $\lambda_{i}$ is the mean free path of interactions when particle propagates inside Galaxy, $\tau_{i}$ is the decay time of particle of type $i$, and $\gamma$ is particle Lorentz factor. In this study, we employ a publicly available version of the DRAGON code to propagate CRs from sources to Earth Evoli et al. (2017). The maximum value of Galactocentric radius is $R_{G}=10\rm~{}kpc$ and halo size is $H_{G}=4\rm~{}kpc$. The spatial distribution of microquasars is unknown. For microquasars like SS 433 or Sgr V4641, it follows the distribution of massive stars, thus have similar spatial distribution with Galactic SNRs. For the source distribution, we adopt the Ferriere2001 model Ferriere (2001), based on the population studies on SNRs. The spatial distribution of the diffusion coefficient is constant and parametrized as $D=23.8\times 10^{28}(\mathcal{R}/250\rm~{}GV)^{\delta}{\rm cm}^{2}{\rm~{}s}^{-1}$ with $\delta=0.45$ for $\mathcal{R}<250\rm~{}GV$, while $\delta=1/3$ for $\mathcal{R}\geqslant 250\rm~{}GV$ Murase and Fukugita (2019); Kimura et al. (2018b). The chosen diffusion coefficient is determined by the observed Boron to Carbon (B/C) ratio.
We adjust the energy injection rate of CRs to match the CR data at PeV energies, which requires $Q_{\rm cr,tot}\sim 1.3\times 10^{42}\rm~{}erg~{}kpc^{-3}~{}yr^{-1}$. The energy injection rate integrated over the Galactic volume is $L_{\rm cr,tot}=Q_{\rm CR,tot}V_{G}\simeq 1.5\times 10^{38}\rm~{}erg~{}s^{-1}$, where $V_{G}\sim 2\pi R_{G}^{2}\times H_{G}\sim 1\times 10^{68}\rm~{}cm^{3}$ is the Galactic volume. Our scenario is consistent with the value required to explain CRs in the knee energy range if we have around $3-10$ microquasars in our Galaxy. For our scenario to work, microquasars need to have extended jets of $\sim 100$ pc. Only two objects (SS433 and V4641 Sgr) exhibit such a feature. The other 3 microquasars, GRS 1915, Cyg X-1, MAXI J1820, do not have such large-scale extended jets, although they have more compact jets. Since we can observe only less than half of our Galaxy, 3-10 microquasars with extended jets sounds more reasonable. Note that the required number of microquasars with extended jets depend on the injection energy given by Eq. II.2.
In Fig. 3, we present the predicted energy spectrum and mean mass $\langle{\rm ln}(A)\rangle$ of CRs from Galactic microquasars in our shear acceleration scenario. Note the shaded region in the lower panel about $\langle{\rm ln}(A)\rangle$ are affected by low- and high-energy components. In our scenario, the CR composition ratio is determined by the Galactic CR composition ratio at the rigidity of $\mathcal{R}_{\rm inj}=E_{\rm inj}/Z$. Since $E_{\rm inj}$ is uncertain, we examine several values of $E_{\rm inj}$ (see the next paragraph). In Fig. 3, we adopt the CR composition at 300 GV from AMS-02 observation  Aguilar et al. (2018a, b, 2015, 2021, 2020). Our model demonstrates that CRs from microquasars could explain both proton and all-particle spectrum around knee. The hard proton spectrum below the peak energy (at a few PeV) and a steep power-law decay beyond the peak predicted by our model agrees well with the LHAASO observations. Furthermore, the model self-consistently predicts the spectra of helium, intermediate-mass, and heavy nuclei around knee, with their relative abundance normalized to the local Galactic CR abundance at 100 GV. Future precision measurements of light and heavy nuclei will provide a critical test of this scenario.
For comparison, we also adopt the CR composition inferred from AMS-02 observations at reference rigidities of 100 GV and 1 TV Aguilar et al. (2018a, b, 2015, 2021, 2020). We also consider a scenario where injected CRs follow solar abundances, $f_{\odot}$, as expected if they are swept from the interstellar medium (ISM) Lodders (2010). The adopted composition ratios are summarized in Table 2. As demonstrated in Ref. Caprioli et al. (2017), the nonthermal tail of nuclei is enhanced by a factor of $(A/Z)^{2}$ for efficient diffusive acceleration in nonrelativistic shocks. Our analysis shows that reaccelerating low-energy CRs (100 GV – 1 TV) yields a mean logarithmic mass $\langle\ln A\rangle$ roughly consistent with observations near the CR knee. In contrast, models assuming solar or enhanced solar compositions fail to reproduce the observed mass distribution. These results might indicate that the reacceleration of pre-existing low-energy CRs provides a plausible explanation for the composition trends observed in the knee region.
Z A $f_{A,\rm 300GV}$ $f_{A,\rm 100GV}$ $f_{A,\rm 1TV}$ $f_{\odot}$ $f_{\odot}^{\rm enhance}$
P 1 1 0.787 0.803 0.767 0.922 0.747
He 2 4 0.194 0.18 0.212 0.077 0.250
CNO 7 16 0.0145 0.013 0.0157 0.007 0.003
NeMgAlSi 12 28 0.00374 0.00352 0.004 0.001 0.0006
Fe 26 56 0.00121 0.00097 0.0014 0.00002 0.0001
Table 2: Injected CR composition at the rigidity 300 GV, 100 GV, and 1 TV for different composition models. In addition, we also include solar composition $f_{\odot}$ and enhanced solar composition $f_{\odot}^{\rm enhance}$ by a factor of $(A/Z)^{2}$.
IV Connection to low- and high-energy components
In this section, we aim to demonstrate that our microquasar model can be combined with other astrophysical models explaining the CR spectrum at different wavelengths. In particular, we provide a possible model, in which the whole CR spectrum from around GeV to 100 EeV can be explained with four main sources, including supernova remnants (SNR, Gal pop1), microquasar (MQ, Gal pop2), clusters of galaxies (CG, ExGal pop1) and radio galaxies (RQ, ExGal pop2). Note that the main purpose here is to demonstrate that the reacceleration model works not only in the large-scale jet-cocoon system of microquasars but also in radio galaxies.
For CRs from Galactic SNRs, we calculate the observed energy spectrum of each element with DRAGON code, and the normalization is adjusted to match the AMS-02 data. We adopt an broken-power law injection spectrum of CRs from SNRs to fit the data. See Appendix for the details. The escaped CR spectrum typically have exponential cutoff beyond the maximum acceleration energy. We set the maximum rigidity from Galactic SNRs to $\mathcal{R}_{\rm max}=0.4\rm~{}PV$.
Figure 4: An overall picture of the CR spectrum from Galactic to extragalactic components. Gal pop1: CRs from supernova remnants (SNR) Reynolds (2008); Blasi (2013) (dotted curve); Gal pop2: CRs from microquasars (MQ) (dashed curve); ExGal pop1: CRs from extragalactic sources, e.g., cluster of galaxies Murase et al. (2008) (dotted-dashed curve); ExGal pop2: UHECRs from radio galaxies (RG) Kimura et al. (2018a) (long dashed curve). The data points are the same as those in Fig. 3, with the AMS data are taken from Refs. Aguilar et al. (2018a, b, 2015, 2021, 2020), and the mean mass distribution measured by Auger is taken from Ref. Yushkov (2020).
For CRs in the knee energy range, we use the CR spectrum that comes directly from Monte Carlo simulations, same as Fig. 3. At energies above around 100 PeV to EeV, CRs from extragalactic sources begin to dominate. At this energy range, the second knee appears, where the spectrum is further steepened from $\gamma\sim 3$ to $\gamma\sim 3.3$, because of the limitation of the maximum energy of the second Galactic component based on the rigidity-dependent modeling. The physical origin of the CRs beyond the second knee up to the ankle has been proposed by various models, and it has been shown that an additional component so-called the B-component is required Gaisser et al. (2016). This component may be responsible for the proton subankle (hardening) feature around 100 PeV observed by KASCADE-Grande Apel et al. (2013) , and the proposed models include the cluster of galaxies Murase et al. (2008, 2013); Simeon et al. (2025), Galactic wind Thoudam et al. (2016); Mukhopadhyay et al. (2023), Wolf-Rayet star explosions Thoudam et al. (2016). In addition, super-knee CR accelerators such as Galactic neutron star merger remnants Kimura et al. (2018b) and interacting supernovae Ptuskin and Zirakashvili (2003, 2005) may explain another hardening feature around $10^{16.22}$ eV reported by TALE Abbasi et al. (2018).
For the subankle proton component, we consider cluster of galaxies for the demonstrative purpose. We consider the injected CRs have composition similar to solar composition, but enhanced by a factor of $(A/Z)^{2}$, see the last column in Tabel 2 about $f_{\odot}^{\rm enhance}$. Note that our results are unaffected by the composition assumption unless the CRs are dominated by heavy nuclei, and such nuclei can be largely disintegrated inside the clusters Kotera et al. (2009); Fang and Murase (2018). The source spectral index above 10-100 PeV is $s_{\rm cr}=2.4$ and the maximum acceleration rigidity $\mathcal{R}_{\rm max}=3.7\times 10^{17}\rm~{}eV$, following Ref. Murase et al. (2008). We propagate CRs from an ensemble of extragalactic sources that are uniformly distributed with the characteristic distance between sources $d_{s}=32.5\rm~{}Mpc$. The number density of galaxy clusters is around $10^{-6}\rm~{}Mpc^{-3}$, which leads to $d_{s}\sim 100$ Mpc. The observed flux is obtained with Eq. 6 in Ref. González et al. (2021). We assume a uniform extragalactic magnetic field $B_{\rm rms}=1\rm~{}nG$ and coherence length $L_{\rm coh}=1\rm~{}Mpc$.
UHECRs above EeV energies are produced from radio galaxies via the shear-espresso acceleration mechanism Caprioli (2015); Kimura et al. (2018a); Mbarek and Caprioli (2019); Mbarek et al. (2023); Seo et al. (2024); Wang et al. (2024); Mbarek et al. (2025). Here, we adopt the predicted spectrum and composition model of UHECRs in KMZ18, where CRs with rigidity $\sim 1$ TV in radio galaxies are re-accelerated by shear in kpc-scale jets. Since radio galaxies are more evolved than our Galaxy, KMZ18 considers that the metal composition is enhanced by a ratio of 3, compared to our Galaxy. Here, we further enhance the iron composition ratio by a factor of 3 to improve the fit to the latest UHECR data. The further enhancement of iron composition ratio in radio galaxies is reasonable because elliptical galaxies have lower star-formation rates than our Galaxy. In radio galaxies, core-collapse supernovae do not frequently occur because of their low star-formation rates, while type Ia supernovae occur because explosions of white dwarfs require significant time lag after star-formation activities. Then, type Ia supernovae inject iron into the ISM, enhancing the iron composition ratio. In addition, we consider a larger value of magnetic field strength in the cocoon with $B_{\rm coc}=4.5\rm~{}\mu G$, where the peak energy of the injection spectrum shown in Fig. 3 in KMZ18 is increased to $\mathcal{R}_{\rm max}\approx 2.4\rm~{}EV$
We found that the overall CR spectrum and composition could be explained by the above four populations. However, one should keep in mind that uncertainty is currently large, and this demonstrates that the reacceleration of Galactic CRs plays a key role in reshaping the mass distribution of accelerated CRs at different energy ranges.
V Summary and Discussion
In this work, we studied the reacceleration and escape of CRs in microquasars within the framework of the shear acceleration mechanism. Our results showed that the CRs from microquasars could explain the observed proton energy spectrum above the PeV energy range. The predicted all-particle spectrum and composition are consistent with the LHAASO observations, where the relative fraction of difference nuclei is constrained by direct measurements around rigidity from 10 GV to 1 TV. The maximum acceleration energy of CRs from shear acceleration in microquasars is close to the knee around 3 PeV, and the escaping CRs have a hard spectrum.
In our scenario, the magnetic field strength in jets and cocoons should be as high as $B_{\rm jet}\sim 50-100\rm~{}\mu G$ and $B_{\rm coc}\sim 10-30\rm~{}\mu G$ to achieve the PeV CR acceleration. Both field strengths are somewhat higher than expected, but the energetics arguments are not violated. The magnetic luminosity in the jet can be estimated to be
$L_{B}$ $=\frac{B_{\rm jet}^{2}}{4\pi}r_{\rm jet}^{2}\beta_{\rm jet}c$ (7)
$\simeq 3.2\times 10^{38}\left(\frac{B_{\rm jet}}{\rm 50\mu G}\right)^{2}\left(\frac{r_{\rm jet}}{\rm 3~{}pc}\right)\left(\frac{\beta_{\rm jet}}{0.3}\right)\rm~{}erg~{}s^{-1}.$
This is slightly lower than the Eddington luminosity, $L_{\rm Edd}\sim 10^{39}(M/10~{}M_{\odot})\rm~{}erg~{}s^{-1}$. The required cocoon magnetic field strength, $B_{\rm coc}\sim 10\rm~{}\mu G$, is a few times higher than the typical ISM value, $B_{\rm ISM}\sim 3\rm~{}\mu G$, and thus, the shock compression of the ISM would be able to naturally achieve the required $B_{\rm coc}$.
Given the universality of the shear acceleration mechanism, we consider a unified model in which microquasars dominate observed Galactic CRs around knee, while radio galaxies are responsible for the observed UHECRs. We found that the whole CR range from GeV to 100 EeV could be explained given that SNRs supply most low-energy Galactic CRs and another population (e.g., galaxy clusters) could contribute at subankle energies. Our model could be tested with future CR experiments with improved measurements of the composition and anisotropy.
B.T.Z. is supported in China by National Key R&D program of China under the grant 2024YFA1611402. This work is partly supported by KAKENHI Nos. 22K14028, 21H04487, 23H04899 (SSK), and 20H05852 (KM). SSK acknowledges the support by the Tohoku Initiative for Fostering Global Researchers for Interdisciplinary Sciences (TI-FRIS) of MEXT’s Strategic Professional Development Program for Young Researchers. Numerical computations were in part carried out on Cray XD2000 at the Center for Computational Astrophysics, National Astronomical Observatory of Japan. K.M. is supported by the NSF Grant Nos. AST-2108466, AST-2108467 and AST-2308021. KM also acknowledges the SUGAR 2024 workshop in Madison, Wisconsin, which led to the idea of shear acceleration.
Appendix A Dependence on source parameters
Here, we present the results for parameters that are different from the default one in Table 1, as indicated in Table 3. The predicted energy spectrum of escaping CRs is summarized in Fig. 5. The predicted CR intensity on Earth is presented in Fig. 6 and Fig. 7, respectively. We can see the peak energy predicted with Model-B is slightly below PeV, which is unable to explain the observed proton knee structure.
Model $r_{\rm jet}$ $B_{\rm jet}$ $\beta_{\rm jet}$ $r_{\rm coc}$ $B_{\rm coc}$ $l_{\rm coh}$ $l_{\rm jet}$
[pc] [$\rm\mu G$] [pc] [$\rm\mu G$] [pc] [pc]
A (default) 4 50 0.2 60 20 1.8 100
B 3 50 0.3 60 10 1.8 100
C 3 100 0.4 60 30 12 100
Table 3: Model parameters used in Monte Carlo simulations.
Figure 5: The energy spectrum of CRs predicted in the shear acceleration mechanism from Monte Carlo simulations. Three different parameter sets are considered.
Figure 6: Similar to Fig. 3, but with parameters from Model-B.
Figure 7: Similar to Fig. 3, but with parameters from Model-C.
Appendix B Propagate CRs in the Galaxy
In Fig. 8, we show the observed energy spectrum of CR species calculated from DRAGON code. The injection spectrum from Galactic SNRs is assumed to have broken powerlaw formula,
$\frac{dN}{d\mathcal{R}}\propto\begin{cases}\mathcal{R}^{-\gamma_{1}}{\rm exp}(-\mathcal{R}/\mathcal{R}_{\rm max}),&\mathcal{R}\leq\mathcal{R}_{12}\\ \mathcal{R}^{-\gamma_{2}}{\rm exp}(-\mathcal{R}/\mathcal{R}_{\rm max}),&\mathcal{R}_{12}<\mathcal{R}\leq\mathcal{R}_{23}\\ \mathcal{R}^{-\gamma_{3}}{\rm exp}(-\mathcal{R}/\mathcal{R}_{\rm max}),&\mathcal{R}_{23}<\mathcal{R}\leq\mathcal{R}_{34}\\ \mathcal{R}^{-\gamma_{4}}{\rm exp}(-\mathcal{R}/\mathcal{R}_{\rm max}),&\mathcal{R}_{34}<\mathcal{R}\leq\mathcal{R}_{45}\\ \mathcal{R}^{-\gamma_{5}}{\rm exp}(-\mathcal{R}/\mathcal{R}_{\rm max}),&\mathcal{R}>\mathcal{R}_{45}\end{cases}$ (8)
where $\mathcal{R}_{\rm max}=0.4\rm~{}PV$. The adopted parameters are listed in  4 and the energy spectrum is presented in Fig. 8.
H He C N O Ne
$\gamma_{1}$ 1.8 2 2 1.9 2 1.9
$\gamma_{2}$ 2.4 2.28 2.35 2.4 2.38 2.4
$\gamma_{3}$ 2.26 2.16 2.15 2.24 2.15 2.24
$\gamma_{4}$ 2.16 2.1 2.15 2.24 2.15 2.24
$\gamma_{5}$ 2.4 2.28 2.35 2.4 2.38 2.4
$\mathcal{R}_{12}$ 7 7 7 7 7 7
$\mathcal{R}_{23}$ 335 165 165 330 165 330
$\mathcal{R}_{34}$ 1000 1000 1000 1000 1000 1000
$\mathcal{R}_{45}$ 5000 5000 5000 5000 5000 5000
Table 4: Parameters used in DRAGON for the propagation of CRs from Galactic SNRs. Note the parameters for nuclei heavier than Neon is keep the same, where we not listed here.
Figure 8: Observed energy spectrum of single element CRs from Galactic SNRs.